OpenStudy (anonymous):

when the expression x^3+mx^2+nx-5 is divided by x+2 and x+3, the remainder are 2 and 13 respectively. Find the values m and n

OpenStudy (anonymous):

@Coolsector pls help again :)

OpenStudy (anonymous):

so it is the same problem p(x) = x^3+mx^2+nx-5 so p(-2) = 2 p(-3) = 13

OpenStudy (anonymous):

but i really want you to understand why and how to write those two : p(-2) = 2 p(-3) = 13

OpenStudy (anonymous):

p(-2) because x+2=0 transpose x=-2 p(-3) because x+3=0 x=-3

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

ive got the answers 4m-2=15 and 9m-3n=45 is that correct?

OpenStudy (anonymous):

i guess you meant 4m-2n=15

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so yes :)

OpenStudy (anonymous):

value of m is 6? pls check

OpenStudy (anonymous):

i dont know, you can find n now and then check them both together with the initial equations : p(-2) = 2 p(-3) = 13